Phoebe is trying new diet. She gets 35 grams of fat a week to "spend" on her two favorite foods, French fries and Ice cream. A serving of French fries contains 3 grams of fat and a serving of Ice cream contains 2 gram of fat. Phoebe's MU ice cream is 25-I( where I is the number of serving of Ice cream she has consumed. Phoebe's MU French fries=30-4F( F= servings of fries she's consumed). To maximize her utility, how many servings of French fries and how many servings of ice cream should she consume?
1
Expert's answer
2015-10-23T10:10:30-0400
P - an amount of fat in one product
l, f - amount of products, that she consumes
To find an answer we should solve a system of equations:
1. MUl /Pl=MUf /Pf - condition that maximize total utility
2. Pl*l+Pf*f=I - budget constraint

1. (25-l)/2=(30-4*f)/3
2. 2*l+3*f=35
from this system of equations l=13 f=3.